# On the Performance of Dual-Antenna Repeater Assisted Bi-Static MIMO ISAC

### v2

@@ Abstract
This paper presents a framework for target detection and downlink data transmission in a repeater-assisted bi-static integrated sensing and communication system. A repeater is an active scatterer that retransmits incoming signals with a complex gain almost instantaneously, thereby enhancing sensing performance by amplifying the echoes reflected by the targets. The same mechanism can also improve downlink communication by mitigating coverage holes. However, the repeater introduces noise and increases interference at the sensing receiver, while also amplifying the interference from target detection signals at the downlink users. The proposed framework accounts for these sensing-communication trade-offs and demonstrates the potential benefits achievable through a carefully designed precoder at the transmitting base station. In particular, our finding is that a higher value of probability of detection can be attained with considerably lower target radar-cross-section variance by deploying repeaters in the target hot-spot areas.
@@ I Introduction
Repeaters (also known as full-duplex relays) are active scatterers, capable of receiving and retransmitting signals instantly with amplification [Jianan_TWC]. They can be deployed on a large scale to enhance coverage for multiple-input multiple-output (MIMO) wireless systems [Erik_WCL, Sara_Vieira_Erik_Mag, Jianan_TWC, Byoung_Magazine, Sergiy_WCL, Emil_repeater]. Reference [Sara_Vieira_Erik_Mag] demonstrated that repeaters can procure a spectral-efficiency (SE) similar to that of distributed MIMO, while obviating the front-haul signalling overhead. Furthermore, [RIS_Makki_IEEE_Acess] indicated that repeaters can outperform reconfigurable intelligent surface (RIS)-assisted systems, with the additional advantage of repeaters being band/frequency selective. The attractive implementation aspects of repeaters have led to their inclusion in the 5G NR standards since 3rd Generation Partnership Project (3GPP) Release $18$ [3gpp].
@@
On parallel avenue, integrated sensing and communication (ISAC) offers efficient spectral and hardware utilization, and is actively being pushed forward as part of 3GPP Release $19$, a 5G-advanced/pre-6G initiative [ISAC_6G]. A detailed survey of ISAC is beyond the scope of this paper; readers can refer to [ISAC_survey, Zhang_Survey_2022, Michail_TWC] and references therein, and specifically for bi-static ISAC[Wang_TWC, Xiao_TIT]. The authors in [Erik_EUSIPCO] considered repeater-assisted mono-static ISAC, [^1] In a mono-static setting, the base station (BS) is required to be full-duplex, which in turn incurs additional signal processing and hardware overhead for self-interference cancellation. where the repeater was used to improve the coverage for downlink communication users. In contrast to [Erik_EUSIPCO], we investigate the feasibility of bi-static ISAC in the presence of a repeater (specifically, a dual-antenna repeater as in [Sara_Vieira_Erik_Mag, Fig. $2$]), focusing on the effects of the repeater on the performance of both target detection and downlink communication.
@@
In bi-static ISAC, one of the BSs transmits precoded downlink data and a dedicated signal for target detection. The other BS performs joint clutter-channel and radar cross-section (RCS) estimation, as well as target detection, using the reflected echoes. In this setting, ideally, a repeater can enhance the overall ISAC performance in two ways: (a) by amplifying the echo reflected from the target, which in turn can improve the detection performance; (b) by removing coverage-holes for the downlink users.
@@
However, the repeater also amplifies communication-to-sensing and sensing-to-communication interference. Specifically, the amplified sensing signal can incur increased interference for the communication users. Contrarily, for sensing, should the amplified downlink signals (directly via the repeater) overwhelm the reflected amplified echo of the target, detection performance at the sensing receiver can severely degrade. To this end, this paper investigates these tradeoffs and discusses the choices of precoders at the transmit BS that can mitigate this interference while retaining the ISAC performance improvement due to the repeater. Moreover, the developed generalized likelihood ratio test (GLRT) framework is agnostic of the channel models and applicable to diverse propagation conditions among the repeater, BS, and users.
@@ II System Model
We consider a bi-static ISAC system where the transmit BS, with $N_{{\mathsf{t}}}$ antennas, beamforms a precoded sensing signal towards a target hot-spot area while simultaneously serving $K$ downlink single-antenna users. The system also contains a dual-antenna repeater. The receiver BS, with $N_{{\mathsf{t}}}$ antennas, performs target detection based on the reflected echoes from the target, also via the repeater. The system model from the communication point of view is illustrated in Fig. 1(a) and the sensing point of view in Fig. 1(b). The system entails several desired and interference channels, whose definitions are:
  • ${\mathbf{f}}_{n}{\in\mathbb{C}^{{N_{{\mathsf{t}}}}}}$: downlink channel from the transmit BS to the $n$th user.
@@
  • ${\mathbf{C}}{\in\mathbb{C}^{{N_{{\mathsf{r}}}\times N_{{\mathsf{t}}}}}}$: the composite clutter channel. Specifically, the clutter channel consisting of the LoS and NLoS paths caused by permanent obstacles can be measured beforehand and canceled [Behdad_Ozlem_TWC]. The unknown part corresponds to the NLoS paths caused by temporary obstacles, which are encased in $\mathbf{C}$}.
Apart from the clutter channel, we assume perfect channel state information (CSI)at the receiver for target detection as a necessary first step to understand the overall system performance with an intuitive and tractable formulation. [^2] The composite channel via the target is $\alpha\boldsymbol{a}_{{\mathsf{r}}}\boldsymbol{a}_{{\mathsf{t}}}^{T}$. We assume only $\boldsymbol{a}_{{\mathsf{r}}}\boldsymbol{a}_{{\mathsf{t}}}^{T}$ to be known, as the BS beamforms to a specific sensing zone whose associated angle of arrival (AoA) and angle of departure (AoD) are pre-calibrated. However, $\alpha$ is unknown and will be estimated as a part of the GLRT later. Further, let $\nu\in\mathbb{C}$ be the repeater amplification factor. The target is characterized by the associated RCS, denoted by $\alpha$, which follows $\alpha{\sim\mathcal{CN}\left({0,\sigma_{{\mathtt{T}}}^{2}}\right)}$, $\sigma_{{\mathtt{T}}}^{2}$ being the RCS variance. The RCS is unknown at the receiver and needs to be estimated. Finally, we adopt a model wherein target detection is performed in a given hotspot, and a larger area can be covered by sending sensing signals to such hotspot areas serially [Behdad_Ozlem_TWC].
@@ III Signalling Scheme
where, $\rho$ is the total transmit power at the BS, $\pi_{n}$ and $\pi_{{\mathtt{T}}}$ denote the fraction of the total power allocated to the $n$th downlink user and for sensing, respectively. Also, ${\mathbf{p}}_{n}$ and ${\mathbf{p}}_{{\mathtt{T}}}$ are the corresponding unit-normalized precoding vectors for the $n$th downlink user and the target. Similarly, $s_{n}\in\mathbb{C}$ and $s_{{\mathtt{T}}}\in\mathbb{C}$ are unit-energy, statistically independent symbols. Also, ${\mathbb{E}\left[{\left\|{\mathbf{x}}[\tau]\right\|_{2}^{2}}\right]}\leq\rho$, satisfying the total transmit power budget. Finally, $\tau_{{\mathtt{L}}}$ is the total duration over which sensing signals are transmitted, which can also be a coherence block. Given this, we can write the input-output receive ($y_{{\mathtt{R}}}^{{\mathtt{I}}}$) and transmit ($y_{{\mathtt{R}}}^{{\mathtt{O}}}$) signals at the repeater terminals as:
@@ III-1 Sensing with repeater
The signal received at the BS can be expressed as:
${\mathbf{y}}_{\mathtt{BS}}[\tau]=\alpha\boldsymbol{a}_{{\mathsf{r}}}\boldsymbol{a}_{{\mathsf{t}}}^{T}{\mathbf{x}}[\tau]+\boldsymbol{b}_{{\mathsf{r}}}y_{{\mathtt{R}}}^{{\mathtt{O}}}[\tau]+{\mathbf{G}}_{{\mathtt{B}}}{\mathbf{x}}[\tau]+{\mathbf{C}}{\mathbf{x}}[\tau]+{\mathbf{w}}_{{\mathtt{B}}}[\tau],$ (3)
@@
where the equivalent sensing channel ${\mathbf{R}}[\tau]$, equivalent clutter channel $\dot{{\mathbf{C}}}$, and the overall sensing noise $\dot{{\mathbf{w}}}_{{\mathtt{B}}}[\tau]$, respectively, can be written as:
@@
III-2 Downlink communication with repeater
For downlink communication analysis, we drop the explicit dependence of index $\tau$. The signal received at the $n$th downlink user can be expressed as
@@ III-2 Downlink communication with repeater
Note that the term $\nu\alpha h_{n}g{\left\{\boldsymbol{a}_{{\mathsf{t}}}^{T}{\mathbf{x}}\right\}}$" in (6) can be subsumed in the effective downlink channel of the users, i.e., $\left\{{\mathbf{f}}_{n}\right\}$}; and is therefore not explicitly shown in (7). This is because the user is agnostic of the presence of the target, and any reflection from it can be treated as from any other environmental scatterers.
Next, based on (4) and (7), we present GLRT-based target detection framework and downlink signal-to-interference-plus-noise ratio (SINR) analysis.
@@ IV RCS Estimation & Target Detection
and $\lambda_{{\mathsf{s}}}$ is the threshold used by the detector, which is selected to achieve a desired false alarm probability. The final result of the GLRT is given in the following lemma.
@@
See Appendix -A. ∎
The complexity of the GLRT is dominated by a matrix inversion of dimension $(N_{{\mathsf{t}}}N_{{\mathsf{r}}}+1)\times(N_{{\mathsf{t}}}N_{{\mathsf{r}}}+1)$, and hence is $\mathcal{O}\left((N_{{\mathsf{t}}}N_{{\mathsf{r}}}+1)^{3}\right)$. Next, only the parameter remains to be computed in the sensing noise covariance $\boldsymbol{\Sigma}_{{\mathsf{s}}}[\tau]={\mathbb{E}\left[\dot{{\mathbf{w}}}_{{\mathtt{B}}}[\tau]\dot{{\mathbf{w}}}_{{\mathtt{B}}}^{H}[\tau]\right]}$, which is evaluated (treating transmitted symbols as pilots) as
$\boldsymbol{\Sigma}_{{\mathsf{s}}}[\tau]={\mathbb{E}\left[\tilde{{\mathbf{G}}}_{{\mathtt{B}}}{\mathbf{x}}[\tau]{\mathbf{x}}^{H}[\tau]\tilde{{\mathbf{G}}}_{{\mathtt{B}}}^{H}\right]}+{\left\{{\big\lvert{\nu}\big\rvert}^{2}\sigma^{2}_{\mathtt{R}}\right\}}\boldsymbol{b}_{{\mathsf{r}}}\boldsymbol{b}_{{\mathsf{r}}}^{H}+\sigma^{2}_{\mathtt{BS}}{\mathbf{I}}_{N_{{\mathsf{r}}}}.$
Next, if we assume that elements of $\tilde{{\mathbf{G}}}_{{\mathtt{B}}}$ are independent and identically distributed (i.i.d.) $\mathcal{CN}\left({0,\zeta^{2}}\right)$}, $\zeta^{2}$ capturing the power of the residual inter-BS interference [AC_CRM_TCOM_2024], then $\mathbb{E}\left[\tilde{{\mathbf{G}}}_{{\mathtt{B}}}{\mathbf{x}}[\tau]{\mathbf{x}}^{H}[\tau]\tilde{{\mathbf{G}}}_{{\mathtt{B}}}^{H}\right]$} evaluates to $\zeta^{2}{\left\|{\mathbf{x}}[\tau]\right\|_{2}^{2}}{\mathbf{I}}_{N_{{\mathsf{r}}}}$, which completes the calculation of $\boldsymbol{\Sigma}_{{\mathsf{s}}}[\tau]$.
@@
We note that a noisy repeater can severely degrade the sensing performance, which is captured by ${\big\lvert{\nu}\big\rvert}^{2}\sigma^{2}_{\mathtt{R}}$ in the sensing noise covariance. The effect of the amplified downlink signals (directly via the repeater) is implicitly captured in the clutter covariance matrix in (11c).
V Downlink SE Analysis
In downlink, based on (1) and (7), the signal received at the $n$th user, with $\dot{{\mathbf{f}}}_{n}\triangleq{{\mathbf{f}}_{n}+\nu h_{n}\boldsymbol{b}_{{\mathsf{t}}}}$, can be expressed as:
$y_{\mathtt{UE},n}=$ $\sqrt{\rho}\sqrt{\pi_{n}}\dot{{\mathbf{f}}}_{n}^{T}{\mathbf{p}}_{n}s_{n}+\sqrt{\rho}\sum\limits_{n^{\prime}=1,n^{\prime}\neq n}^{K}\sqrt{\pi_{n^{\prime}}}\dot{{\mathbf{f}}}_{n}^{T}{\mathbf{p}}_{n^{\prime}}s_{n^{\prime}}$
@@ V Downlink SE Analysis
where the second and the third term correspond to multi-user interference and interference due to a dedicated signal for target detection. The instantaneous SINR at the $n$th downlink user can then be written as
${\mathtt{SINR}}_{n}=\frac{\rho\pi_{n}{\big\lvert{\dot{{\mathbf{f}}}_{n}^{T}{\mathbf{p}}_{n}}\big\rvert}^{2}}{\rho\sum\limits_{n^{\prime}=1,n^{\prime}\neq n}^{K}\pi_{n^{\prime}}{\big\lvert{\dot{{\mathbf{f}}}_{n}^{T}{\mathbf{p}}_{n^{\prime}}}\big\rvert}^{2}+\rho\pi_{{\mathtt{T}}}{\big\lvert{\dot{{\mathbf{f}}}_{n}^{T}{\mathbf{p}}_{{\mathtt{T}}}}\big\rvert}^{2}+{\mathbb{E}\left[{\big\lvert{\dot{w}_{\mathtt{UE}}}\big\rvert}^{2}\right]}},$
where the noise variance can be computed as
${\mathbb{E}\left[{\big\lvert{\dot{w}_{\mathtt{UE}}}\big\rvert}^{2}\right]}={\big\lvert{\nu}\big\rvert}^{2}{\big\lvert{h_{n}}\big\rvert}^{2}\sigma^{2}_{\mathtt{R}}+\sigma^{2}_{\mathtt{UE}}.$ (17)
Next, for the choice of the precoders, the transmitter can either consider the composite channel $\left\{\dot{{\mathbf{f}}}_{n}\right\}$} or be completely agnostic of the repeater and consider $\left\{{{\mathbf{f}}}_{n}\right\}$}. We consider regularized zero-forcing (ZF) precoding scheme for the communication user, in which case ${\mathbf{p}}_{n}$ becomes ${\mathbf{p}}_{n}=\epsilon_{n}{\left(\sum\limits_{n^{\prime}=1}^{K}\dot{{\mathbf{f}}}_{n^{\prime}}\dot{{\mathbf{f}}}_{n^{\prime}}^{H}+\zeta_{{\mathtt{ZF}}}{\mathbf{I}}_{N_{{\mathsf{t}}}}\right)}^{-1}\dot{{\mathbf{f}}}_{n},$ where $\epsilon_{n}$ is the normalizer to ensure ${\left\|{\mathbf{p}}_{n}\right\|_{2}^{2}}=1$ and $\zeta_{{\mathtt{ZF}}}>0$ is a regularizer. For ${\mathbf{p}}_{{\mathtt{T}}}$ we can consider one of the following: Target centric: In this case, the transmitter beamforms towards the target, i.e., ${\mathbf{p}}_{{\mathtt{T}}}=\epsilon_{{\mathtt{T}}}\dot{{\mathbf{f}}}_{{\mathtt{T}}}$, with $\dot{{\mathbf{f}}}_{{\mathtt{T}}}=\boldsymbol{a}_{{\mathsf{t}}}$ and $\epsilon_{{\mathtt{T}}}$ is an appropriate normalizer. Communication centric: We null the (potentially destructive) interference from the sensing signal to the users by projecting ${\mathbf{p}}_{{\mathtt{T}}}$ onto the nullspace of the subspace spanned by the users’ channel vectors. In this case, ${\mathbf{p}}_{{\mathtt{T}}}=\epsilon_{{\mathtt{T}}}{\left({\mathbf{I}}_{N_{{\mathsf{t}}}}-{\mathbf{U}}{\mathbf{U}}^{H}\right)}\dot{{\mathbf{f}}}_{{\mathtt{T}}}$, where $\mathbf{U}$} is a unitary matrix formed with the orthonormal columns that span the column space of $\left\{\dot{{\mathbf{f}}}_{1},\dot{{\mathbf{f}}}_{2},\ldots,\dot{{\mathbf{f}}}_{K}\right\}$}.

### v3

@@ Abstract
This paper presents a framework for target detection and downlink (DL) data transmission in a repeater-assisted bi-static integrated sensing and communication system. A repeater is an active scatterer that retransmits incoming signals with a complex gain almost instantaneously, thereby enhancing sensing performance by amplifying the echoes reflected by the targets. The same mechanism can also improve DL communication by mitigating coverage holes. However, the repeater introduces noise and increases interference at the sensing receiver, while also amplifying the interference from target detection signals at the DL users. The proposed framework accounts for these sensing-communication trade-offs and demonstrates the potential benefits achievable through a carefully designed precoder at the transmitting base station. In particular, our finding is that a higher value of probability of detection can be attained with considerably lower target radar-cross-section variance by deploying repeaters in the target hot-spot areas.
@@ I Introduction
Repeaters (also known as full-duplex relays) are active scatterers, capable of receiving and retransmitting signals instantly with amplification [1]. They can be deployed on a large scale to enhance coverage for multiple-input multiple-output (MIMO) wireless systems [2, 3, 1, 4, 5, 6]. Reference [3] demonstrated that repeaters can procure a spectral-efficiency (SE) similar to that of distributed MIMO, while obviating the front-haul signalling overhead. Furthermore, [7] indicated that repeaters can outperform reconfigurable intelligent surface (RIS)-assisted systems, with the additional advantage of repeaters being band/frequency selective. The attractive implementation aspects of repeaters have led to their inclusion in the 5G NR standards since 3rd Generation Partnership Project (3GPP) Release $18$ [8].
@@
On parallel avenue, integrated sensing and communication (ISAC) offers efficient spectral and hardware utilization, and is actively being pushed forward as part of 3GPP Release $19$, a 5G-advanced/pre-6G initiative [9]. A detailed survey of ISAC is beyond the scope of this paper; readers can refer to [10, 11, 12] and references therein, and specifically for bi-static ISAC[13, 14, 15]. Specifically, [15] optimized RIS-assisted bistatic sensing. The authors in [16] considered repeater-assisted mono-static ISAC, [^1] In a mono-static setting, the base station (BS) is required to be full-duplex, which in turn incurs additional signal processing and hardware overhead for self-interference cancellation. where the repeater was used to improve the coverage for DL communication users. In contrast to [16], we investigate the feasibility of bi-static ISAC in the presence of a repeater (specifically, a dual-antenna repeater [3, Fig. $2$]), focusing on the effects of the repeater on the performance of both target detection and DL communication.
@@
In bi-static ISAC, one of the BSs transmits precoded DL data and a dedicated signal for target detection. The other BS performs joint clutter-channel and radar cross-section (RCS) estimation, and target detection, using the reflected echoes. In this setting, ideally, a repeater can enhance ISAC performance in two ways: (a) by amplifying the echo reflected from the target, which in turn can improve the detection performance; (b) by removing coverage-holes for the DL users.
@@
However, the repeater also amplifies communication-to-sensing and sensing-to-communication interference. Specifically, the amplified sensing signal can incur increased interference for the communication users. Contrarily, for sensing, should the amplified DL signals (directly via the repeater) overwhelm the reflected amplified echo of the target, detection performance at the sensing receiver can severely degrade. To this end, this paper investigates these tradeoffs and discusses the choices of precoders at the transmit BS that can mitigate this interference while retaining the ISAC performance improvement due to the repeater. Moreover, the developed generalized likelihood ratio test (GLRT) framework is agnostic of the channel models and applicable to diverse propagation conditions among the repeater, BS, and users.
@@ II System Model
We consider a bi-static ISAC system where the transmit BS, with $N_{{\mathsf{t}}}$ antennas, beamforms a precoded sensing signal towards a target hot-spot area while simultaneously serving $K$ DL single-antenna users. The system also contains a dual-antenna repeater. The receiver BS, with $N_{{\mathsf{t}}}$ antennas, performs target detection based on the reflected echoes from the target, also via the repeater. The system model from the communication point of view is illustrated in Fig. 1(a) and the sensing point of view in Fig. 1(b). The system entails several desired and interference channels, whose definitions are:
  • ${\mathbf{f}}_{n}{\in\mathbb{C}^{{N_{{\mathsf{t}}}}}}$: DL channel from the transmit BS to the $n$th user.
@@
  • ${\mathbf{C}}{\in\mathbb{C}^{{N_{{\mathsf{r}}}\times N_{{\mathsf{t}}}}}}$: the composite clutter channel. Specifically, the clutter channel consisting of the LoS and NLoS paths caused by permanent obstacles can be measured beforehand and canceled [17]. The unknown part corresponds to the NLoS paths caused by temporary obstacles, encased in $\mathbf{C}$}.
Apart from the clutter channel, we assume perfect channel state information (CSI)at the receiver for target detection as a necessary first step to understand the overall system performance with an intuitive and tractable formulation. [^2] The composite channel via the target is $\alpha\boldsymbol{a}_{{\mathsf{r}}}\boldsymbol{a}_{{\mathsf{t}}}^{T}$. We assume only $\boldsymbol{a}_{{\mathsf{r}}}\boldsymbol{a}_{{\mathsf{t}}}^{T}$ to be known, as the BS beamforms to a specific sensing zone whose associated angle of arrival (AoA) and angle of departure (AoD) are pre-calibrated. However, $\alpha$ is unknown and will be estimated as a part of the GLRT later. Further, let $\nu\in\mathbb{C}$ be the repeater amplification factor. The target is characterized by the associated RCS, denoted by $\alpha$, which follows $\alpha{\sim\mathcal{CN}\left({0,\sigma_{{\mathtt{T}}}^{2}}\right)}$, $\sigma_{{\mathtt{T}}}^{2}$ being the RCS variance. The RCS is unknown at the receiver and needs to be estimated. Finally, we adopt a model wherein target detection is performed in a given hotspot, and a larger area can be covered by sending sensing signals to such hotspot areas serially [17].
@@ III Signalling Scheme
where, $\rho$ is the total transmit power at the BS, $\pi_{n}$ and $\pi_{{\mathtt{T}}}$ denote the fraction of the total power allocated to the $n$th DL user and for sensing, respectively. Also, ${\mathbf{p}}_{n}$ and ${\mathbf{p}}_{{\mathtt{T}}}$ are the corresponding unit-normalized precoding vectors for the $n$th DL user and the target. Similarly, $s_{n}\in\mathbb{C}$ and $s_{{\mathtt{T}}}\in\mathbb{C}$ are unit-energy, statistically independent symbols. Also, ${\mathbb{E}\left[{\left\|{\mathbf{x}}[\tau]\right\|_{2}^{2}}\right]}\leq\rho$, satisfying the total transmit power budget. Finally, $\tau_{{\mathtt{L}}}$ is the total duration over which sensing signals are transmitted, which can also be a coherence block. Given this, we can write the input-output receive ($y_{{\mathtt{R}}}^{{\mathtt{I}}}$) and transmit ($y_{{\mathtt{R}}}^{{\mathtt{O}}}$) signals at the repeater terminals as:
@@ III-1 Sensing with repeater
The signal received at the BS is
${\mathbf{y}}_{\mathtt{BS}}[\tau]=\alpha\boldsymbol{a}_{{\mathsf{r}}}\boldsymbol{a}_{{\mathsf{t}}}^{T}{\mathbf{x}}[\tau]+\boldsymbol{b}_{{\mathsf{r}}}y_{{\mathtt{R}}}^{{\mathtt{O}}}[\tau]+{\mathbf{G}}_{{\mathtt{B}}}{\mathbf{x}}[\tau]+{\mathbf{C}}{\mathbf{x}}[\tau]+{\mathbf{w}}_{{\mathtt{B}}}[\tau],$
@@
where the equivalent sensing channel ${\mathbf{R}}[\tau]$, equivalent clutter channel $\dot{{\mathbf{C}}}$, and the overall sensing noise $\dot{{\mathbf{w}}}_{{\mathtt{B}}}[\tau]$, are:
@@
III-2 DL communication with repeater
For DL communication analysis, we drop the explicit dependence of index $\tau$. The signal received at the $n$th DL user can be expressed as
@@ III-2 DL communication with repeater
Note that the term $\nu\alpha h_{n}g{\left\{\boldsymbol{a}_{{\mathsf{t}}}^{T}{\mathbf{x}}\right\}}$" in (5) can be subsumed in the effective DL channel of the users, i.e., $\left\{{\mathbf{f}}_{n}\right\}$}; and is therefore not explicitly shown in (6). This is because the user is agnostic of the presence of the target, and any reflection from it can be treated as from any other environmental scatterers.
@@ IV RCS Estimation & Target Detection
and $\lambda_{{\mathsf{s}}}$ is the threshold used by the detector, which is selected to achieve a desired false alarm probability.
@@
See Appendix Appendix A: Proof of Lemma 1. ∎
The complexity of the GLRT is dominated by a matrix inversion of dimension $(N_{{\mathsf{t}}}N_{{\mathsf{r}}}+1)\times(N_{{\mathsf{t}}}N_{{\mathsf{r}}}+1)$, and hence is $\mathcal{O}\left((N_{{\mathsf{t}}}N_{{\mathsf{r}}}+1)^{3}\right)$. Next, only the parameter remains to be computed in the sensing noise covariance $\boldsymbol{\Sigma}_{{\mathsf{s}}}[\tau]={\mathbb{E}\left[\dot{{\mathbf{w}}}_{{\mathtt{B}}}[\tau]\dot{{\mathbf{w}}}_{{\mathtt{B}}}^{H}[\tau]\right]}$, which is evaluated (treating transmitted symbols as pilots) as $\boldsymbol{\Sigma}_{{\mathsf{s}}}[\tau]={\mathbb{E}\left[\tilde{{\mathbf{G}}}_{{\mathtt{B}}}{\mathbf{x}}[\tau]{\mathbf{x}}^{H}[\tau]\tilde{{\mathbf{G}}}_{{\mathtt{B}}}^{H}\right]}+{\left\{{\big\lvert{\nu}\big\rvert}^{2}\sigma^{2}_{\mathtt{R}}\right\}}\boldsymbol{b}_{{\mathsf{r}}}\boldsymbol{b}_{{\mathsf{r}}}^{H}+\sigma^{2}_{\mathtt{BS}}{\mathbf{I}}_{N_{{\mathsf{r}}}}.$ Next, if we assume that elements of $\tilde{{\mathbf{G}}}_{{\mathtt{B}}}$ are independent and identically distributed (i.i.d.) $\mathcal{CN}\left({0,\zeta^{2}}\right)$}, $\zeta^{2}$ capturing the power of the residual inter-BS interference [18], then $\mathbb{E}\left[\tilde{{\mathbf{G}}}_{{\mathtt{B}}}{\mathbf{x}}[\tau]{\mathbf{x}}^{H}[\tau]\tilde{{\mathbf{G}}}_{{\mathtt{B}}}^{H}\right]$} evaluates to $\zeta^{2}{\left\|{\mathbf{x}}[\tau]\right\|_{2}^{2}}{\mathbf{I}}_{N_{{\mathsf{r}}}}$, which completes the calculation of $\boldsymbol{\Sigma}_{{\mathsf{s}}}[\tau]$.
@@
We note that a noisy repeater can severely degrade the sensing performance, which is captured by ${\big\lvert{\nu}\big\rvert}^{2}\sigma^{2}_{\mathtt{R}}$ in the sensing noise covariance. The effect of the amplified DL signals (directly via the repeater) is implicitly captured in the clutter covariance matrix in (10c).
V DL SE Analysis
In DL, based on (1) and (6), the signal received at the $n$th user, with $\dot{{\mathbf{f}}}_{n}\triangleq{{\mathbf{f}}_{n}+\nu h_{n}\boldsymbol{b}_{{\mathsf{t}}}}$, can be expressed as:
$y_{\mathtt{UE},n}=$ $\sqrt{\rho}\sqrt{\pi_{n}}\dot{{\mathbf{f}}}_{n}^{T}{\mathbf{p}}_{n}s_{n}+\sqrt{\rho}\sum\nolimits_{n^{\prime}=1,n^{\prime}\neq n}^{K}\sqrt{\pi_{n^{\prime}}}\dot{{\mathbf{f}}}_{n}^{T}{\mathbf{p}}_{n^{\prime}}s_{n^{\prime}}$
@@ V DL SE Analysis
where the second and the third term correspond to multi-user interference and interference due to a dedicated signal for target detection. The instantaneous signal-to-interference-plus-noise ratio (SINR) at the $n$th DL user can then be written as ${\mathtt{SINR}}_{n}=\frac{\rho\pi_{n}{\big\lvert{\dot{{\mathbf{f}}}_{n}^{T}{\mathbf{p}}_{n}}\big\rvert}^{2}}{\rho\sum\limits_{n^{\prime}=1,n^{\prime}\neq n}^{K}\pi_{n^{\prime}}{\big\lvert{\dot{{\mathbf{f}}}_{n}^{T}{\mathbf{p}}_{n^{\prime}}}\big\rvert}^{2}+\rho\pi_{{\mathtt{T}}}{\big\lvert{\dot{{\mathbf{f}}}_{n}^{T}{\mathbf{p}}_{{\mathtt{T}}}}\big\rvert}^{2}+{\mathbb{E}\left[{\big\lvert{\dot{w}_{\mathtt{UE}}}\big\rvert}^{2}\right]}},$ where the noise variance can be computed as ${\mathbb{E}\left[{\big\lvert{\dot{w}_{\mathtt{UE}}}\big\rvert}^{2}\right]}={\left\{{\big\lvert{\nu}\big\rvert}^{2}{\big\lvert{h_{n}}\big\rvert}^{2}\sigma^{2}_{\mathtt{R}}+\sigma^{2}_{\mathtt{UE}}\right\}}.$ Next, for the choice of the precoders, the transmitter can either consider the composite channel $\left\{\dot{{\mathbf{f}}}_{n}\right\}$} or be completely agnostic of the repeater and consider $\left\{{{\mathbf{f}}}_{n}\right\}$}. We consider regularized zero-forcing (ZF) precoding scheme for the communication user, in which case ${\mathbf{p}}_{n}$ becomes ${\mathbf{p}}_{n}=\epsilon_{n}{\left(\sum\nolimits_{n^{\prime}=1}^{K}\dot{{\mathbf{f}}}_{n^{\prime}}\dot{{\mathbf{f}}}_{n^{\prime}}^{H}+\zeta_{{\mathtt{ZF}}}{\mathbf{I}}_{N_{{\mathsf{t}}}}\right)}^{-1}\dot{{\mathbf{f}}}_{n},$ where $\epsilon_{n}$ is the normalizer to ensure ${\left\|{\mathbf{p}}_{n}\right\|_{2}^{2}}=1$ and $\zeta_{{\mathtt{ZF}}}>0$ is a regularizer. For ${\mathbf{p}}_{{\mathtt{T}}}$ we can consider one of the following: Target centric: In this case, the transmitter beamforms towards the target, i.e., ${\mathbf{p}}_{{\mathtt{T}}}=\epsilon_{{\mathtt{T}}}\dot{{\mathbf{f}}}_{{\mathtt{T}}}$, with $\dot{{\mathbf{f}}}_{{\mathtt{T}}}=\boldsymbol{a}_{{\mathsf{t}}}$ and $\epsilon_{{\mathtt{T}}}$ is an appropriate normalizer. Communication centric: We null the (potentially destructive) interference from the sensing signal to the users by projecting ${\mathbf{p}}_{{\mathtt{T}}}$ onto the nullspace of the subspace spanned by the users’ channel vectors. In this case, ${\mathbf{p}}_{{\mathtt{T}}}=\epsilon_{{\mathtt{T}}}{\left({\mathbf{I}}_{N_{{\mathsf{t}}}}-{\mathbf{U}}{\mathbf{U}}^{H}\right)}\dot{{\mathbf{f}}}_{{\mathtt{T}}}$, where $\mathbf{U}$} is a unitary matrix formed with the orthonormal columns that span the column space of $\left\{\dot{{\mathbf{f}}}_{1},\dot{{\mathbf{f}}}_{2},\ldots,\dot{{\mathbf{f}}}_{K}\right\}$}.